OpenStudy (anonymous):

Shelby purchased a plot of land for $27,500. The land appreciates about 2.9% each year. What is the value of the land after nine years? $21,101.13 $35,568.95 $34,677.50 $28,103.94

OpenStudy (johnweldon1993):

so we have the formula V = Vi (1 + r)^t Vi is initial value t is time r is rate (in decimal form) can you plug your numbers into this?

OpenStudy (johnweldon1993):

so lets take it step by step what is the initial value?

OpenStudy (johnweldon1993):

you might be busy....or having a mental block with this....so I'll do it out step by step so you may reference it for future use V = Vi (1 + r)^t Vi is initial value = 27,500 t is time = 9 years r is rate (in decimal form) .029 so V = 27,500 (1 + .029)^9 this equates to 35568.95

OpenStudy (anonymous):

thank you so much, I was having amental block wth this actually, taking me step by step helped me break free from it, thanks :)

OpenStudy (johnweldon1993):

anytime! glad I could help :)